OpenStudy (anonymous):

if the area of the rectangle is 24in^2, find the demensions.

OpenStudy (anonymous):

it has x on the side and x+5 on top if that helps.

OpenStudy (solomonzelman):

Oh, then it makes sense.

OpenStudy (solomonzelman):

\[A= base \times hieght\] so\[A= (x+5) \times x\] we know that the area is equal to 24, so\[24= (x+5)x\] can you solve for x?

OpenStudy (anonymous):

that would become x^2 + 5=24 then I would subtract 5 from 24 making it x^2= 19 right?

OpenStudy (anonymous):

oh wait the 5 would become 5x wouldn't it?

OpenStudy (solomonzelman):

Not exactly, it is x^2+5x=24

OpenStudy (anonymous):

now i'm lost

OpenStudy (anonymous):

would it be 3in x 8in?

OpenStudy (cggurumanjunath):

x^2 + 5x=24 x^2-5x-24=0

OpenStudy (cggurumanjunath):

@becca_baby

OpenStudy (solomonzelman):

@becca_baby, you got it!

OpenStudy (cggurumanjunath):

can u solve for x ?

OpenStudy (cggurumanjunath):

do u know how to factorise x^2-5x-24 @becca_baby ?

OpenStudy (solomonzelman):

Which methods do you know? Quadratic formula Factoring Completing the square (Sorry, CGGURUMANJUNATH, I am just reasking your question in a better way.)

OpenStudy (anonymous):

i understand the quadratic formula and factoring. I am doing math online and my teacher hasn't been of much help

OpenStudy (cggurumanjunath):

i was just continuing from where u had left previously @SolomonZelman

OpenStudy (solomonzelman):

Yeah, I understand, I just got tagged 5 times, so please take over, @CGGURUMANJUNATH.

OpenStudy (cggurumanjunath):

no let us continue together @SolomonZelman

OpenStudy (anonymous):

can I have help with polynomials possibly?

OpenStudy (cggurumanjunath):